# High-Speed Event Vision-Based Tactile Roller Sensor for Large Surface Measurements

### v1

@@ III-A Mechanical Design and Fabrication
As illustrated in Figure 2a, our proposed sensor consists of: a clear, transparent acrylic cylinder as a support structure; a translucent elastomer with a reflective coating; a neuromorphic camera (DVXplorer mini); a white LED ring with diffuser; shafts that attach to the camera holder; bearings to keep the camera holder stationary during rolling; and sealing covers to attach the cylinder and shafts to the outer housing as well as seal the light from the LEDs within the sensor.
@@
(a) Refer to caption 1 Side Connector  2 Side Cover  3 Bearing  4 Rotating Shaft  5 Top Connector  6 Camera Holder  7 DVXplorer mini camera  8 LED Ring with Diffuser  9 Acrylic Cylinder and Elastomer
(b) Refer to caption
@@
Figure 2: Mechanical Design of Roller: (a) Exploded view of roller components; (b) Front view (left) and side view (right) of assembled roller.
@@
The supporting parts, which were used to connect all the components, consisted of shafts, a camera holder, and sealing covers. The other components (shafts, camera holder, sealing covers) were 3D printed with PLA. Lastly, brass inserts and screws were used to secure the components together. The final assembled roller is shown in Figure 2b.
III-B Event Camera Principles
Unlike conventional cameras that capture intensity information across all pixels at fixed frame rates, event cameras employ a fundamentally distinct, bio-inspired sensing paradigm [28]. Each pixel operates independently and asynchronously, responding dynamically to changes in illumination. Specifically, a pixel monitors the logarithm of the incident light intensity, $L(x,y,t)=\log(I(x,y,t))$. An event $e_{k}=(x_{k},y_{k},t_{k},p_{k})$ is generated at coordinates $(x_{k},y_{k})$ only when the change in this logarithmic intensity, $\Delta L$, relative to its value at the time of the last event from that pixel, surpasses a predefined contrast threshold, ±$\pm C$. The generated event includes a high-resolution timestamp $t_{k}$ and a polarity $p_{k}\in\{-1,+1\}$, signifying the direction of the brightness change (decrease or increase, respectively). Following event generation, the pixel’s internal reference brightness level is updated, priming it for subsequent intensity variations. The output is thus not a sequence of frames, but a sparse, continuous stream of asynchronous events reflecting dynamic scene changes. The unique characteristics of event cameras offer substantial benefits for high-speed tactile inspection. Their inherent high temporal resolution, with events timestamped often at microsecond precision, allows for tactile sensing at high speeds without the motion blur experienced by other roller tactile sensors at similar speeds. Furthermore, event-based cameras benefit from low latency and high dynamic range allowing for potential reduction in computational load and robustness to varying illumination within the sensor housing respectively.
III-C Surface Reconstruction
III-C1 Event-Based Multi-View Stereo (EMVS)
@@
In motion parallax, objects closer to the observer (or camera) appear to move faster across the field of view than objects farther away. This effect is due to the relative difference in angular displacement caused by motion, and it’s a key depth cue used by both biological and artificial vision systems to perceive 3D structure from motion. To estimate the 3D structure of the scanned surface from the event stream generated by the moving roller, we adapt the multi-view stereo (MVS) space-sweep method, originally developed for frame-based cameras [50] and later extended to event cameras (EMVS) [51]. This technique leverages the known viewpoints (poses) of the event camera over time. The core idea involves discretizing a volume of interest, $\bar{D}\in\mathbb{R}^{w\times h\times N_{z}}$, encompassing the potential 3D scene, where $w$, $h$, and $N_{z}$ represent the dimensions of the volume. Each captured event, $e_{k}=(x_{k},y_{k},t_{k},p_{k})$, is back-projected through this volume as a 3D ray using the camera’s calibrated intrinsic parameters and its known pose at time $t_{k}$.
@@
A Disparity Space Image (DSI) is employed to accumulate evidence for surfaces within the volume. The DSI acts as a 3D histogram where each voxel counts the number of back-projected event rays passing through it, typically referenced to a specific viewpoint. Voxels corresponding to real surfaces in the scene are expected to lie at the intersection of rays from multiple viewpoints, resulting in high ray counts. Local maxima in the ray density within the DSI are then identified as potential 3D points belonging to the scanned surface [51, 52]. Processing these maxima using Adaptive Gaussian Thresholding, yields a semi-dense depth map relative to the reference viewpoint.
III-C2 Multi-Reference Depth Map Generation and Fusion
The standard EMVS approach generates a depth map referenced to a single key viewpoint, often chosen as the midpoint $t_{m}=(t_{s}+t_{e})/2$ within a processing time window $[t_{s},t_{e}]$. We select a time window of 20 ms, balancing the need for sufficient event integration for reconstruction accuracy with the requirement for high-speed operation (up to 0.5 m/s) without introducing significant motion artifacts in the DSI.
However, roller-based tactile sensing can suffer from errors due to the sensor’s curved nature, leading to inconsistent contact depth and potential fabrication variations [53]. High-speed motion can also introduce occlusion and warping artifacts. To enhance robustness against these effects and specifically mitigate the depth inconsistency and curvature-induced errors inherent in cylindrical tactile sensing, we extend the EMVS framework by generating and fusing depth maps from multiple reference times within the same time window. Inspired by works demonstrating the benefits of multi-reference objectives in event-based vision [54] and fusion techniques in roller tactile sensing [53], our approach specifically computes depth maps $Z_{s}$, $Z_{m}$, and $Z_{e}$ referenced at the start ($t_{s}$), midpoint ($t_{m}$), and end ($t_{e}$) times of the window, respectively. As demonstrated quantitatively and qualitatively in the Results section (Section V, Figure 9), this fusion strategy significantly improves depth consistency and reduces depth map curvature distortion compared to EMVS.
We propose fusing these individual depth maps using a Bayesian Model Averaging (BMA) strategy (Figure 3), which has proven effective for combining multiple noisy estimates in vision tasks [55]. The fused depth map, $Z_{f}$, is computed as a weighted average of the individual depth maps after warping them to a common reference frame (chosen as the midpoint view, $t_{m}$):
$Z_{f}=\sum_{i\in\{s,m,e\}}w_{i}\cdot W_{i}(Z_{i})$ (1)
@@ III-C2 Multi-Reference Depth Map Generation and Fusion
where $Z_{i}$ is the depth map generated at reference time $t_{i}$, $w_{i}$ is the corresponding fusion weight, and $W_{i}$ is the warping function. The warping function $W_{i}$ relies on standard camera projection models. Event cameras utilize the same optical principles as traditional cameras, allowing the use of the pinhole projection model to relate 3D world points to 2D image points. A 3D point $\mathbf{P}_{W}=[X_{W},Y_{W},Z_{W}]^{T}$ in the world frame is projected onto the image plane as a 2D pixel $\mathbf{p}=[u,v,1]^{T}$ according to:
$z\mathbf{p}=\mathbf{K}[\mathbf{R}|\mathbf{t}]\begin{bmatrix}\mathbf{P}_{W}\\ 1\end{bmatrix}=\mathbf{K}\begin{bmatrix}\mathbf{P}_{C}\\ 1\end{bmatrix}$ (2)
@@
where $z$ is the depth $Z_{C}$ of the point in the camera frame, $\mathbf{K}$ is the $\times 3$ 3 camera intrinsic matrix, $[\mathbf{R}|\mathbf{t}]$ represents the 3×43$\times$ 43 × 4 extrinsic matrix (rotation $\mathbf{R}$ and translation $\mathbf{t}$) transforming world coordinates to camera coordinates $\mathbf{P}_{C}=[X_{C},Y_{C},Z_{C}]^{T}$. The absolute pose of the camera frame relative to the robot base frame, ${}^{B}\mathbf{T}_{C}(t)$, which defines $\mathbf{R}$ and $\mathbf{t}$ relative to the base frame at time $t$, is obtained by combining the robot’s end-effector pose ${}^{B}\mathbf{T}_{E}(t)$ (derived from forward kinematics) with the fixed hand-eye transformation ${}^{E}\mathbf{T}_{C}$, determined via hand-eye geometric calibration [52]. The intrinsic matrix $\mathbf{K}$ is also obtained through calibration using E-Calib, a calibration toolbox for event cameras [56].
@@
    $\begin{bmatrix}\mathbf{P}_{m}\\ 1\end{bmatrix}={}^{m}\mathbf{T}_{i}\begin{bmatrix}\mathbf{P}_{i}\\ 1\end{bmatrix}$ (4)
@@
The original depth value $Z_{i}(u,v)$ is then associated with the calculated warped pixel location $(u^{\prime},v^{\prime})$ in the reference frame $m$ before the weighted averaging (Eq. 1) is performed.
Figure 3: Multi-Reference Fusion Strategy: For each reference time, each incoming ray from a back-projected event votes for the voxels it passes through in the corresponding DSI ray counter. Depth map is extracted from each DSI and these depth maps are fused using BMA.
III-D Parameter Calibration
@@
To establish the ground truth depth map $Z_{gt}$ for the sphere contact, we adapt the methodology used in [49]. When the calibration sphere is pressed onto the sensor, we first identify the resulting circular contact patch in the depth map (visualized in Figure 4). From this, we determine the radius of the pressed circle, $r$ (in pixels), and the coordinates of its center, $(x_{b},y_{b})$ (in pixels). The depth of the sphere’s center, $z_{b}$ (in pixels), can be calculated geometrically as:
$z_{b}=-\sqrt{(R_{sphere})^{2}-r^{2}}$ (6)
For any pixel $(x_{i},y_{i})$ within the identified contact circle, its depth $z_{i}$ (in pixels) must satisfy the sphere equation:
$(x_{i}-x_{b})^{2}+(y_{i}-y_{b})^{2}+(z_{i}-z_{b})^{2}=(R_{sphere})^{2}$ (7)
Solving Equation 7 for $z_{i}$ for all pixels inside the contact circle yields the ground truth depth map in pixel units. This map is then converted to millimeters to obtain $Z_{gt}$ by multiplying with the scale factor $f$. We determined this scale factor $f=0.0422$ mm/pixel by imaging features of known dimensions using a Vernier caliper (varying space between jaws from 2 to 5 mm and averaging results).
@@
III-E Cross-Validation and Weighted Fusion Analysis
[…]
We then evaluated how well each set of parameters $\theta_{i}$ performed on average across all $M=3$ test objects. For each candidate parameter set $\theta_{i}$ (calibrated on object $i$), we calculated its average MAE, $e_{i}$, across all test objects $j$ as follows:
$e_{i}=\frac{1}{M}\sum_{j=1}^{M}\bigl{\|}Z_{f}^{(j)}(\theta_{i})-Z_{gt}^{(j)}\bigr{\|}_{1}$ (9)
where $Z_{f}^{(j)}(\theta_{i})$ is the fused depth map generated for object $j$ using parameters $\theta_{i}$. We then compared these average error values to determine which calibration object yielded the parameters with the best overall generalization performance (i.e., the minimum $e_{i}$).
The results are summarized in Table II. While the object-specific optimal weights vary slightly, the weights calibrated using the sphere yielded the lowest overall average MAE ($e_{\text{Sphere}}=0.0549$ mm) when applied universally, confirming its suitability as a general calibration target.
TABLE II: Cross-validation of BMA fusion weights
Calibrated on $w_{s}$ $w_{m}$ $w_{e}$ Avg. MAE (mm)
Sphere 0.445 0.333 0.222 0.0549
Stairs 0.5 0.333 0.167 0.0652
Sparse Sphere 0.4 0.4 0.2 0.0649
A consistent and perhaps counter-intuitive trend observed across all calibrations is that the start reference view ($t_{s}$) receives the highest (or tied for highest) weight, while the end view ($t_{e}$) consistently receives the lowest. Intuitively, one might expect the midpoint view ($t_{m}$), representing a temporal average, to be most reliable. However, this weighting suggests the information captured early in the time window is most critical for accurate reconstruction in our setup. Several factors related to the EMVS algorithm and rolling contact dynamics may contribute:
  • Pose Confidence: The pose at $t_{s}$ is derived directly from the robot state at the window’s beginning. The pose at $t_{m}$ relies on interpolation and assumes constant velocity within the 20ms window, therefore there could be a disparity between the interpolated pose and actual pose at $t_{m}$.The estimated pose at $t_{e}$ may carry slightly greater uncertainty than at $t_{s}$, stemming from both potential inaccuracies in interpolating the robot pose between controller updates to match precise event timestamps, and the manifestation of small kinematic errors or unmodeled dynamic effects during the 20ms motion interval. Thus, the geometric reference at $t_{s}$ might be inherently the most accurate relative to the events occurring near it.
  • Information Content: As the roller moves, tactile features first appear and generate events with high contrast near the start of the window ($t_{s}$). This information might provide the most stable features for the EMVS algorithm before potential disocclusions occur. Conversely, features rolling out of view near $t_{e}$ might contribute less reliable data or introduce artifacts.
  • Error Accumulation in DSI: The DSI integrates events over the entire window. Errors in pose estimation or event timing later in the window could slightly degrade the DSI’s quality relative to the $t_{e}$ viewpoint compared to its quality relative to the $t_{s}$ viewpoint.
Overall, the results strongly suggest that the initial phase of the time window provides the most pertinent information for robust reconstruction via EMVS for this high-speed tactile rolling application.
III-F Braille Feature Recognition
[…]

### v2

@@ IV-A Mechanical Design and Fabrication
As illustrated in Figure 2, our proposed sensor consists of: a clear, transparent acrylic cylinder as a support structure; a translucent elastomer with a reflective coating; a neuromorphic camera (DVXplorer mini); a white LED ring with diffuser; shafts that attach to the camera holder; bearings to keep the camera holder stationary during rolling; and sealing covers to attach the cylinder and shafts to the outer housing as well as seal the light from the LEDs within the sensor.
@@
Refer to caption 1 Side Connector    2 Side Cover    3 Bearing    4 Rotating Shaft    5 Top Connector    6 Camera Holder    7 DVXplorer mini camera    8 LED Ring with Diffuser    9 Acrylic Cylinder and Elastomer
@@
Figure 2: Mechanical Design of Roller: Exploded view of roller components with roller dimensions and insert of assembled view.
@@
The supporting parts, which were used to connect all the components, consisted of shafts, a camera holder, and sealing covers. The other components (shafts, camera holder, sealing covers) were 3D printed with PLA. Lastly, brass inserts and screws were used to secure the components together. The final assembled roller is shown (as an insert) in Figure 2. The design files for the roller are available at: https://anonymous.4open.science/r/TSMR_design_files-6269/
IV-B Surface Reconstruction
IV-B1 Event-Based Multi-View Stereo (EMVS)
@@
In motion parallax, objects closer to the observer (or camera) appear to move faster across the field of view than objects farther away. This effect is due to the relative difference in angular displacement caused by motion, and it’s a key depth cue used by both biological and artificial vision systems to perceive 3D structure from motion. To estimate the 3D structure of the scanned surface from the event stream generated by the moving roller, we adapt the multi-view stereo (MVS) space-sweep method, originally developed for frame-based cameras [collins_1996_a] and later extended to event cameras (EMVS) [rebecq_2017_emvs]. This technique leverages the known viewpoints (poses) of the event camera over time. The core idea involves discretizing a volume of interest, $\bar{D}\in\mathbb{R}^{w\times h\times N_{z}}$, encompassing the potential 3D scene, where $w$, $h$, and $N_{z}$ represent the dimensions of the volume. Each captured event, $e_{k}=(x_{k},y_{k},t_{k},p_{k})$, is back-projected through this volume as a 3D ray using the camera’s calibrated intrinsic parameters and its known pose at time $t_{k}$. For our purposes, the physical design of the roller sensor provides a strong prior on the depth range at which events occur. We exploit this to optimize the discretization of the volume $\bar{D}$ for both accuracy and computational efficiency. The parameters $w$ and $h$ are set to the camera’s resolution of 640x480 respectively. Furthermore, the depth dimension is constrained to a 4 mm range, from a minimum depth of 43 mm to a maximum depth of 47 mm, corresponding to the expected range of elastomer deformation from the sensor’s 47 mm radius. This volume is discretized into $N_{z}=$ 500 depth planes, yielding a high depth resolution of 8 µm per plane. This constrained search space is critical for achieving both high-speed processing and high-precision reconstruction.
@@
A Disparity Space Image (DSI) is employed to accumulate evidence for surfaces within the volume. The DSI acts as a 3D histogram where each voxel counts the number of back-projected event rays passing through it, typically referenced to a specific viewpoint. Voxels corresponding to real surfaces in the scene are expected to lie at the intersection of rays from multiple viewpoints, resulting in high ray counts. Local maxima in the ray density within the DSI are then identified as potential 3D points belonging to the scanned surface [rebecq_2017_emvs, ayyad_2023_neuromorphic]. Processing these maxima using Adaptive Gaussian Thresholding, yields a semi-dense depth map relative to the reference viewpoint.
Our implementation, in line with the original EMVS framework [rebecq_2017_emvs], conceptually associates a unique pose with every event. In our hardware setup (refer to Section V), the sensor pose is derived from the UR10 robot’s end-effector state, which is logged at 100 Hz. To obtain a pose for any event’s specific microsecond-resolution timestamp, we perform linear interpolation between the two nearest pose readings. For computational efficiency, as is done in EMVS, events are grouped into small temporal packets (comprised of 256 events) and all events in a packet are assigned the interpolated pose of the median event’s timestamp.
@@
The warping function relies on standard camera projection models. Event cameras utilize the same optical principles as traditional cameras, allowing the use of the pinhole projection model to relate 3D world points to 2D image points. A 3D point $\mathbf{P}_{W}=[X_{W},Y_{W},Z_{W}]^{T}$ in the world frame is projected onto the image plane as a 2D pixel $\mathbf{p}=[u,v,1]^{T}$ according to:
$z\mathbf{p}=\mathbf{K}[\mathbf{R}|\mathbf{t}]\begin{bmatrix}\mathbf{P}_{W}\\ 1\end{bmatrix}=\mathbf{K}\mathbf{P}_{C}$ (1)
@@
where $z$ is the depth $Z_{C}$ of the point in the camera frame, $\mathbf{K}$ is the $3\times 3$ camera intrinsic matrix, $\mathbf{R}|\mathbf{t}]$ represents the $3\times 4$ extrinsic matrix (rotation $\mathbf{R}$, which is a $3\times 3$ matrix and translation $\mathbf{t}$, which is $3\times 1$) transforming world coordinates to camera coordinates $\mathbf{P}_{C}=[X_{C},Y_{C},Z_{C}]^{T}$. The absolute pose of the camera frame relative to the robot base frame, $^{B}\mathbf{T}_{C}(t)$, which defines $\mathbf{R}$ and $\mathbf{t}$ relative to the base frame at time $t$, is obtained by combining the robot’s end-effector pose $^{B}\mathbf{T}_{E}(t)$ (derived from forward kinematics) with the fixed hand-eye transformation $^{E}\mathbf{T}_{C}$, determined via hand-eye geometric calibration [ayyad_2023_neuromorphic]. The intrinsic matrix $\mathbf{K}$ is also obtained through calibration using E-Calib, a calibration toolbox for event cameras [salah_2024_ecalib].
@@
    $\begin{bmatrix}\mathbf{P}_{m}\\ 1\end{bmatrix}={}^{m}\mathbf{T}{i}\begin{bmatrix}\mathbf{P}_{i}\\ 1\end{bmatrix}$ (3)
@@
IV-B2 Multi-Reference Depth Map Generation and Fusion
The standard EMVS approach, described above, generates a depth map referenced to a single key viewpoint, often chosen as the midpoint $t_{m}=(t_{s}+t_{e})/2$ within a processing time window $t_{s},t_{e}]$. We select a time window of 20 ms, balancing the need for sufficient event integration for reconstruction accuracy with the requirement for high-speed operation (up to 0.5 m/s) without introducing significant motion artifacts in the DSI. This 20 ms time window refers to the total duration of event data used to generate a single depth map, and this window contains numerous event packets, each with a distinct pose, ensuring multi-view stereo reconstruction.
However, roller-based tactile sensing can suffer from errors due to the sensor’s curved nature, leading to inconsistent contact depth and potential fabrication variations [li_2025_cyclic]. The curved elastomer means that even when contacting a flat surface, the contact points have varying depths relative to the camera. Second, the small contact area means any given surface feature is only observed over a short rolling distance within a short time window, resulting in a small baseline for multi-view stereo, which can make depth estimation more challenging. High-speed motion can also introduce occlusion and warping artifacts. To enhance robustness against these effects and specifically mitigate the depth inconsistency and curvature-induced errors inherent in cylindrical tactile sensing, we extend the EMVS framework by generating and fusing depth maps from multiple reference times within the same time window.
Inspired by works demonstrating the benefits of multi-reference objectives in event-based vision [shiba_2024_secrets] and fusion techniques in roller tactile sensing [li_2025_cyclic], our approach extends the EMVS framework through a three-stage fusion process. First, we generate three separate DSIs using the same 20 ms stream of events and their associated poses. Each DSI is constructed relative to a different reference pose: one at the start ($t_{s}$), one at the midpoint ($t_{m}$), and one at the end ($t_{e}$) of the time window. Although the input data is identical, the differing reference frames result in distinct DSI geometries and consequently, slightly different ray count distributions. Second, as is done in [rebecq_2017_emvs], for each of the three DSIs, we extract a deterministic depth map by performing a maximization along each pixel’s viewing ray to find the depth plane with the highest vote count. This yields three independent depth maps: $Z_{s}$, $Z_{m}$, and $Z_{e}$ referenced at the start ($t_{s}$), midpoint ($t_{m}$), and end ($t_{e}$) times of the window, respectively.
Third, we fuse these individual depth maps using a Bayesian Model Averaging (BMA) strategy (Figure 3), which has proven effective for combining multiple noisy estimates in vision tasks [sun_2018_bayesian]. This process averages out noise and inconsistencies present in the individual reconstructions, mitigating both DSI discretization artifacts and the impact of noise and occlusions. The final fused depth map, $Z_{f}$, is computed as a weighted average of the three deterministic maps after warping them to a common reference frame (chosen as the midpoint view, $t_{m}$):
$Z_{f}=\sum_{i\in\{s,m,e\}}w_{i}\cdot W_{i}(Z_{i})$ (5)
where $Z_{i}$ is the depth map generated at reference time $t_{i}$, $w_{i}$ is the corresponding fusion weight, and $W_{i}$ is the warping function described in Section IV-B1. The original depth value $Z_{i}(u,v)$ is associated with the calculated warped pixel location $(u^{\prime},v^{\prime})$ in the reference frame $m$ before the weighted averaging is performed.
Figure 3: Multi-Reference Fusion Strategy: The roller sensor captures depth maps at three reference times within the time window: start ($t_{s}$), midpoint ($t_{m}$), and end ($t_{e}$). The blue dashed horizontal line indicates the rolling path of the sensor. The light blue triangular frustums represent the DSI ray counters for each reference time. The red arrows represent incoming viewing rays from back-projected events, which vote for all the DSI voxels they pass through. Depth maps are extracted from each DSI and fused using Bayesian Model Averaging (BMA).
IV-C Parameter Calibration
@@
To establish the ground truth depth map $Z_{gt}$ for the sphere contact, we adapt the methodology used in [zhang_gelroller_2024]. For calibration, the sensor is rolled over the calibration sphere while recording event data. The sphere is mounted in a fixed position on an optical breadboard such that the sphere passes directly under the camera’s field of view during rolling, ensuring consistent alignment (refer to setup in Figure 6). Then, the resulting circular contact patch in the depth map is identified (visualized in Figure 4). From this, we determine the radius of the pressed circle, $r$ (in pixels), and the coordinates of its center, $(x_{b},y_{b})$ (in pixels). The physical radius of the calibration sphere, $R_{sphere}$, is known in millimeters. The measured contact radius, $r$, (in pixels) is then converted to millimeters by multiplying by the scale factor $f$ (mm/pixel), where $f=0.0422$ mm/pixel. This scale factor was determined by imaging the gap between the jaws of a Vernier caliper indented directly on the sensor surface at known separations. Two readings were taken at each of four separations (2 mm, 3 mm, 4 mm, and 5 mm), the scale factor was computed for each reading, and the eight resulting values were averaged to obtain the final scale factor $f$. The depth of the sphere’s center, $z_{b}$ (in mm), can then be calculated geometrically as:
$z_{b}=-\sqrt{R_{sphere}^{2}-(r\cdot f)^{2}}$ (6)
For any pixel $(x_{i},y_{i})$ within the identified contact circle, its depth $z_{i}$ (in mm) must satisfy the sphere equation:
$\big((x_{i}-x_{b})\cdot f\big)^{2}+\big((y_{i}-y_{b})\cdot f\big)^{2}+(z_{i}-z_{b})^{2}=R_{sphere}^{2}$ (7)
Here, $(x_{i}-x_{b})$ and $(y_{i}-y_{b})$ are multiplied by the scale factor $f$ to convert from pixels to millimeters before solving the equation. Solving Equation 7 for $z_{i}$ for all pixels inside the contact circle yields the ground truth depth map directly in millimeters.
@@
IV-D Braille Feature Recognition
To demonstrate the sensor’s capability for fine feature detection at high speed, a task complementary to 3D geometric reconstruction, we implemented a Braille character recognition pipeline. This task serves as a well-established benchmark in the tactile sensing community for evaluating high-speed spatial resolution, allowing for a direct comparison with state-of-the-art systems like that of Potdar et al. [potdar_high-speed_2024]. This method specifically evaluates the sensor’s ability to capture fine textural details directly from the event stream, a key aspect of tactile inspection that is distinct from geometric reconstruction. The process begins by applying motion compensation [Salah2023, Gallego_CMAX_2018] to the recorded event stream as the sensor rolls over the Braille plate. Given the constant linear velocity of the roller, $v_{roller}$, which is commanded to and precisely controlled by the UR10 robotic arm as described in the Experimental Setup (Sec. V), we can determine the corresponding optical flow $\mathbf{\dot{U}}=(\dot{u},\dot{v})$ of features moving across the image plane with $[u,v]$ as the pixel coordinates. The roller’s motion involves predominantly horizontal motion, therefore $\dot{v}\approx 0$ and $\dot{u}=\frac{-F}{Z}v_{roller}$, where $F$ is the focal length of the camera and $Z$ is the depth of the scene (assumed to be constant depth of 47 mm which is the radius of the roller with maximum deformation of 4.5 mm).